# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 4.   Characteristic Factors
4.2.   The general case of 4.2
The induction that we will use to prove 4.2 is similar to the PET-induction scheme utilized in [Ber87]. Given a function $f\in\mathcal{H}$ of polynomial growth, we call the smallest $d\in\mathbb{N}$ for which $|f(t)|\ll t^{d}$ the degree of $f$, and denote it by $\deg(f)$. Consider the relation $f\sim_{\operatorname{PET}}g$ if and only if $\deg f=\deg g>\deg(f-g)$ and note that $\sim_{\operatorname{PET}}$ defines an equivalence relation on $\mathcal{H}$.
Now given a finite collection =$\mathcal{F}=\{f_{1},\ldots,f_{k}\}$ of functions of polynomial growth from a Hardy field $\mathcal{H}$, let $d_{\max}$ be a number that is bigger or equal than the degree of any function in $\mathcal{F}$. Also, for each $d\in\{1,\dots,d_{\max}\}$, let $m_{d}$ denote the number of equivalence classes of the set $\{f\in\mathcal{F}:\deg(f)=d\}$ with respect to the equivalence relation $\sim_{\operatorname{PET}}$. The vector $(m_{1},\dots,m_{d_{\max}})$ is called the characteristic vector of $\mathcal{F}$. We order characteristic vectors by letting $(m_{1},\dots,m_{d_{\max}})<(\tilde{m}_{1},\dots,\tilde{m}_{d_{\max}})$ if the maximum $j$ for which $m_{j}\neq\tilde{m}_{j}$ satisfies $m_{j}<\tilde{m}_{j}$.
Proof of 4.2.
We prove the theorem by induction on the characteristic vector of =$\mathcal{F}=\{f_{1},\dots,f_{k}\}$. The case when $d_{\max}=1$ (i.e., all functions in $\mathcal{F}$ have degree $1$) follows from 4.3. We can therefore assume that $f_{k}$ has degree at least $2$ and that the theorem has been proved for all families whose characteristic vector is strictly smaller than that of $\mathcal{F}$.
@@
where the suprema are taken over all functions $h_{1},\ldots,h_{k-1}\in L^{\infty}(X)$ with $\|h_{i}\|_{L^{\infty}}\leqslant 1$ and all $a\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$. As we did in the proof of 4.3 in the previous subsection, instead of asking for the suprema over $h_{1},\ldots,h_{k-1}$ and $a$, it suffices to show that for any $h_{1,N},\ldots,h_{k-1,N}\in L^{\infty}(X)$ with $\|h_{i,N}\|_{L^{\infty}}\leqslant 1$ and any $a_{N}\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$ we have
[…]
where, for convenience, we took $h_{k,N}=h_{k}$ for all $N\in\mathbb{N}$. Then we use 4.6 with $u_{N}(n)=a_{N}(n)\prod_{i=0}^{k}T^{[f_{i}(n)]}h_{i,N}$ and conclude that in order to establish (4.26) it suffices to show that for every $m\in\mathbb{N}$
[…]
Note that $[f_{i}(n+m)]-[f_{1}(n)]=[f_{i}(n+m)-f_{1}(n)]+e_{i,n}$ where $e_{i,n}\in\{-1,0,1\}$ and $[f_{i}(n)]-[f_{1}(n)]=[f_{i}(n)-f_{1}(n)]+\tilde{e}_{i,n}$ where $\tilde{e}_{i,n}\in\{-1,0,1\}$. For each vector $v=(v_{1},\dots,v_{k},\tilde{v}_{1},\dots\tilde{v}_{k})\in\{-1,0,1\}^{2k}$, let $A_{v}$ be the set of $n$’s for which
@@
[…]
@@
[…]
@@
Observe that $\lvert\!|\!|\tilde{h}_{2k,N}|\!|\!\rvert_{s}=\lvert\!|\!|h_{k}|\!|\!\rvert_{s}=0$. Also, since $|f_{i}(t)|\ll|f_{k}(t)|$ for all $i=1,\ldots,k-1$, it follows that $|\tilde{f}_{i,m}(t)|\ll|\tilde{f}_{2k,m}(t)|$ for all $i=1,\ldots,2k-1$. Moreover, since the degree of $f_{k}$ is at least $2$, the degree of $\tilde{f}_{2k,m}$ is at least $1$. Thus (4.30) will follow by induction after we show that the characteristic vector $(\tilde{m}_{1},\dots,\tilde{m}_{\ell})$ of ~=$\tilde{\mathcal{F}}=\{\tilde{f}_{1,m},\ldots,\tilde{f}_{2k,m}\}$ is strictly smaller than the characteristic vector $(m_{1},\dots,m_{\ell})$ of =$\mathcal{F}=\{f_{1},\ldots,f_{k}\}$.
Indeed, for each $f_{i}$ which is not equivalent to $f_{1}$, the functions $\tilde{f}_{i,m}$ and $\tilde{f}_{k+i,m}$ are equivalent to each other. Moreover, if $f_{i}(t)$ and $f_{j}(t)$ are not equivalent to $f_{1}(t)$, then $\tilde{f}_{i,m}$ is equivalent to $\tilde{f}_{j,m}$ if and only if $f_{i}(t)$ is equivalent to $f_{j}(t)$. Letting $d$ be the degree of $f_{1}(t)$, this shows that $\tilde{m}_{j}=m_{j}$ for all $j>d$. Finally, if $f_{i}(t)$ is equivalent to $f_{1}(t)$, then both $f_{i}(t)-f_{1}(t)$ and $f_{i}(t+m)-f_{1}(t)$ have degree smaller than that of $f_{i}(t)$. This shows that $\tilde{m}_{d}<m_{d}$. Therefore $(\tilde{m}_{1},\dots,\tilde{m}_{\ell})$ is strictly smaller than $(m_{1},\dots,m_{\ell})$ and we are done. ∎
5.   Proof of B
In this section we give the proof of B. The first step is to use 4.2, which was proved in the previous section, together with the structure theory of Host-Kra to reduce B to the following special case:
@@ 5.   Proof of B
B holds when $(X,\mathcal{B},\mu,T)$ is an ergodic nilsystem.
A measure preserving system is called an $s$-step pro-nilsystem if it is (isomorphic in the category of measure preserving systems to) an inverse limit of $s$-step nilsystems. In other words, $(X,\mathcal{B},\mu,T)$ is an $s$-step pro-nilsystem if there exist $T$-invariant $\sigma$-algebras ${\mathcal{B}}_{1}\subset{\mathcal{B}_{2}}\subset\cdots\subset{\mathcal{B}}$ such that ${\mathcal{B}}=\bigcup{\mathcal{B}}_{n}$ and for all $n$, the system $(X,{\mathcal{B}}_{n},\mu,T)$ is isomorphic to an $s$-step nilsystem. The following result from [HK05a] gives the relation between the uniformity seminorms and pro-nilsystems.
Theorem 5.2 ([HK05a]).
@@
    a function $h\in L^{\infty}$ is measurable with respect to ${\mathcal{Z}}_{s}$ if and only if it belongs to the set $\left\{h\in L^{\infty}(X):\int_{X}h\cdot h^{\prime}~{}\mathrm{d}\mu=0\quad\forall h^{\prime}\in L^{\infty}(X,{\mathcal{B}},\mu)\text{ with }\lvert\!|\!|h^{\prime}|\!|\!\rvert_{s+1}=0\right\}$.
Proof of B.
Let $\mathcal{H}$ be a Hardy field, let $f_{1},\ldots,f_{k}\in\mathcal{H}$ be of polynomial growth. Let $W\in\mathcal{H}$ be compatible with $f_{1},\dots,f_{k}$, in the sense that Property 1 holds. Let $(X,\mathcal{B},\mu,T)$ be an invertible measure preserving and let $h_{1},\dots,h_{k}\in L^{\infty}(X)$ and $A\in\mathcal{B}$ with $\mu(A)>0$. Using a usual ergodic decomposition argument, we may assume without loss of generality that the system is ergodic.
If $\lim_{t\to\infty}f_{i}(t)$ is finite for some $f_{i}$, then neither condition of A holds so parts 3 and 2 hold vacuously. As every $f_{i}\in\mathcal{H}$ is eventually monotone, this also implies that $[f_{i}(n)]$ is eventually constant, and hence part 1 holds if it holds when $f_{i}$ is removed from $\{f_{1},\dots,f_{k}\}$. Iterating this observation we may assume that $\lim_{t\to\infty}|f_{i}(t)|=\infty$ for every $i$. A similar argument shows that we can assume that $\lim_{t\to\infty}|f_{i}(t)-f_{j}(t)|=\infty$ for every $i\neq j$.
Let $s\in\mathbb{N}$ be given by 4.2 and let ${\mathcal{Z}}_{s}$ be given by 5.2. In view of Theorems 4.2 and 5.2 we can assume that all the functions $h_{i}$ are measurable in ${\mathcal{Z}}_{s}$, which is equivalent to assuming that $(X,\mathcal{B},\mu,T)$ is itself an $s$-step pro-nilsystem. Therefore there exists a sequence of $\sigma$-algebras $\mathcal{B}_{1}\subset\mathcal{B}_{2}\subset\cdots\subset\mathcal{B}$ such that for each $\ell\in\mathbb{N}$ the system $(X,\mathcal{B}_{\ell},\mu,T)$ is an ergodic nilsystem.
It now follows from 5.1 that the conclusions 1 and 2 of B hold after replacing each $h_{i}$ with $\mathbb{E}[h_{i}\mid\mathcal{B}_{\ell}]$, and that
@@
where $H_{\ell}=\mathbb{E}[1_{A}\mid\mathcal{B}_{\ell}]$ and $\mathbb{E}[\cdot\mid\mathcal{B}_{\ell}]$ denotes the conditional expectation on the $\sigma$-subalgebra $\mathcal{B}_{\ell}$. Since $\mathbb{E}[h_{i}\mid\mathcal{B}_{\ell}]\to h_{i}$ and $H_{\ell}\to 1_{A}$ in $L^{p}$ for every $p<\infty$ as $\ell\to\infty$, we conclude that the conclusions of B also hold for $h_{1},\dots,h_{k}$ and $A$. ∎
The rest of this section is devoted to proving 5.1.
Outline of the proof of 5.1:
The first step in the proof of 5.1 is a reduction to the case when the nilmanifold $X$ is of the form $X=G/\Gamma$ for a connected and simply connected nilpotent Lie group $G$. These conditions on $G$ ensure that the nilsystem is embedable into a flow (i.e. a $\mathbb{R}$-action) on the same nilmanifold. This reduction is achieved at the price of losing ergodicity of the $\mathbb{Z}$-action; nevertheless, the $\mathbb{R}$-flow is ergodic. Then another reduction is performed, to avoid local obstructions.
These reductions allow us to reformulate 5.1 as a statement about uniform distribution of certain sequences in a nilmanifold, which we prove by drawing from recent results obtained in [Ric23].
@@ 5.1.   Reducing to connected and simply connected groups
The first step in the proof of 5.1 is essentially a reduction to the case when $X$ is a nilmanifold $G/\Gamma$ where the corresponding nilpotent Lie group $G$ is connected and simply connected. This is a somewhat technical and by now standard reduction but which is crucial in order to deal with the rounding function and to use the equidistribution results from [Ric23]. It turns out that the additional structure of nilsystems allow us to upgrade $L^{2}$ convergence into pointwise convergence. We now state this case of 5.1 as a separate theorem.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property 1. Then for any nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, where $X=G/\Gamma$ and $G$ is connected and simply connected, we have:
  1. [label=(),ref=(),leftmargin=*]
  2. 1.
@@
  3. 2.
    If $f_{1},\ldots,f_{k}$ satisfy condition 1 of A then for any $h_{1},\ldots,h_{k}\in C(X)$ and any $x\in X$
@@
  4. 3.
    If $f_{1},\ldots,f_{k}$ satisfy condition 2 of A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
[…]
Proof of 5.1 assuming 5.3.
Let $(X,\mathcal{B}_{X},\mu_{X},T)$ be an ergodic nilsystem.
In view of 2.1 we can represent $X=G/\Gamma$ as a closed subsystem of another nilsystem $(Z,\mathcal{B}_{Z},\mu_{Z},S)$ where $Z=\hat{G}/\hat{\Gamma}$ for a connected and simply connected nilpotent Lie group $\hat{G}$ which contains $G$ as a closed and rational subgroup. Letting $\tilde{G}$ be the connected component of the identity in $G$ and letting $a\in G$ be such that $T$ corresponds to left multiplication by $a$, we also have that $G=\overline{a^{\mathbb{Z}}\tilde{G}}$ and $\hat{G}=\overline{a^{\mathbb{R}}\tilde{G}}$. If $G=\hat{G}$ then $G$ is connected and simply connected and the desired conclusion follows directly from 5.3. From now on suppose that $G\neq\hat{G}$.
To establish parts 1 and 2 of B for the nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, it suffices to consider continuous functions $h_{i}$, for they are dense in $L^{2}$. By the Tietze-Urysohn extension theorem, any continuous function $h\in C(X)$ can be extended to a continuous function on $Z$. The conclusions now follows directly from parts 1 and 2 of 5.3.
Before proving part 3 of B for the nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, we need some facts about the relation between $X$ and $Z$. For every $y\in X$ the set $D:=\{t\in\mathbb{R}:a^{t}y\in X\}=\{t\in\mathbb{R}:a^{t}\in G\}$ forms a closed subgroup of $\mathbb{R}$. We can’t have $D=\mathbb{R}$ for this would imply $G=\hat{G}$, so $D$ must be discrete. Notice that $D$ is also independent of $y$. Let $c=\min\{t>0:t\in D\}$. Let $R:=S^{c}:X\to X$. Since $1\in D$, it follows that $c=\tfrac{1}{m}$ for some $m\in\mathbb{N}$ and hence $R^{m}=T$
@@
Now, fix $A\subset X$ with $\mu_{X}(A)>0$. Let $B=\bigcup_{t\in[0,c)}S^{t}A$. It follows that $S^{-t}B\cap X=A$ for every $t\in[0,c)$ and hence $\mu_{Z}(B)=m\int_{0}^{c}\mu_{X}(a^{-t}B)~{}\mathrm{d}t=\mu_{X}(A)>0$. In view of part 3 of 5.3 (and the choice of $W$) applied to $(Z,\mathcal{B}_{Z},\mu_{Z},S)$ we deduce that
@@
as long as the functions $f_{i}$ satisfy either condition 2 or 1 of A. Since $X$ is $T$ invariant, for any $n_{1},\dots,n_{k}\in\mathbb{Z}$ and any $t\in[0,c)$ we have

### v3

@@ 4.  Characteristic Factors
4.2.  The general case of ˜4.2
The induction that we will use to prove ˜4.2 is similar to the PET-induction scheme utilized in [Ber87]. Given a function $f\in\mathcal{H}$ of polynomial growth, we call the smallest $d\in\mathbb{N}$ for which $|f(t)|\ll t^{d}$ the degree of $f$, and denote it by $\deg(f)$. Consider the relation $f\sim_{\operatorname{PET}}g$ if and only if $\deg f=\deg g>\deg(f-g)$ and note that $\sim_{\operatorname{PET}}$ defines an equivalence relation on $\mathcal{H}$.
Now given a finite collection $\mathcal{F}=\{f_{1},\ldots,f_{k}\}$ of functions of polynomial growth from a Hardy field $\mathcal{H}$, let $d_{\max}$ be a number that is bigger or equal than the degree of any function in $\mathcal{F}$. Also, for each $d\in\{1,\dots,d_{\max}\}$, let $m_{d}$ denote the number of equivalence classes of the set $\{f\in\mathcal{F}:\deg(f)=d\}$ with respect to the equivalence relation $\sim_{\operatorname{PET}}$. The vector $(m_{1},\dots,m_{d_{\max}})$ is called the characteristic vector of $\mathcal{F}$. We order characteristic vectors by letting $(m_{1},\dots,m_{d_{\max}})<(\tilde{m}_{1},\dots,\tilde{m}_{d_{\max}})$ if the maximum $j$ for which $m_{j}\neq\tilde{m}_{j}$ satisfies $m_{j}<\tilde{m}_{j}$.
Proof of ˜4.2.
We prove the theorem by induction on the characteristic vector of $\mathcal{F}=\{f_{1},\dots,f_{k}\}$. The case when $d_{\max}=1$ (i.e., all functions in $\mathcal{F}$ have degree $1$) follows from ˜4.3. We can therefore assume that $f_{k}$ has degree at least $2$ and that the theorem has been proved for all families whose characteristic vector is strictly smaller than that of $\mathcal{F}$.
@@
where the suprema are taken over all functions $h_{1},\ldots,h_{k-1}\in L^{\infty}(X)$ with $\|h_{i}\|_{L^{\infty}}\leqslant 1$ and all $a\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$. As we did in the proof of ˜4.3 in the previous subsection, instead of asking for the suprema over $h_{1},\ldots,h_{k-1}$ and $a$, it suffices to show that for any $h_{1,N},\ldots,h_{k-1,N}\in L^{\infty}(X)$ with $\|h_{i,N}\|_{L^{\infty}}\leqslant 1$ and any $a_{N}\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$ we have
$\limsup_{N\to\infty}~\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a_{N}(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i,N}\right\|_{L^{2}}=\,0,$ (4.26)
where, for convenience, we took $h_{k,N}=h_{k}$ for all $N\in\mathbb{N}$. Then we use ˜4.6 with $u_{N}(n)=a_{N}(n)\prod_{i=0}^{k}T^{[f_{i}(n)]}h_{i,N}$ and conclude that in order to establish (4.26) it suffices to show that for every $m\in\mathbb{N}$
$\begin{split}\lim_{N\to\infty}\Bigg|\frac{1}{W(N)}&\sum_{n=1}^{N}w(n)a_{N}(n+m)\overline{a}_{N}(n)\\ &\int\prod_{i=1}^{k}T^{[f_{i}(n+m)]}h_{i,N}\cdot T^{[f_{i}(n)]}\overline{h}_{i,N}~\mathrm{d}\mu\Bigg|=\,0.\end{split}$ (4.27)
Note that $f_{i}(n+m)]-[f_{1}(n)]=[f_{i}(n+m)-f_{1}(n)]+e_{i,n}$ where $e_{i,n}\in\{-1,0,1\}$ and $f_{i}(n)]-[f_{1}(n)]=[f_{i}(n)-f_{1}(n)]+\tilde{e}_{i,n}$ where $\tilde{e}_{i,n}\in\{-1,0,1\}$. For each vector $v=(v_{1},\dots,v_{k},\tilde{v}_{1},\dots\tilde{v}_{k})\in\{-1,0,1\}^{2k}$, let $A_{v}$ be the set of $n$’s for which
@@
$\begin{split}\lim_{N\to\infty}\Bigg|\frac{1}{W(N)}&\sum_{n=1}^{N}w(n)a_{N}(n+m)\overline{a}_{N}(n)1_{A_{v}}(n)\\ &\int\prod_{i=1}^{k}T^{[f_{i}(n+m)]}h_{i,N}\cdot T^{[f_{i}(n)]}\overline{h}_{i,N}~\mathrm{d}\mu\Bigg|=\,0.\end{split}$ (4.28)
@@
$\begin{split}\lim_{N\to\infty}&\Bigg|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a_{N}(n+m)\overline{a}_{N}(n)1_{A_{v}}(n)\\ &\int\prod_{i=1}^{k}T^{[f_{i}(n+m)-f_{1}(n)]+v_{i}}h_{i,N}\cdot T^{[f_{i}(n)-f_{1}(n)]+\tilde{v}_{i}}\overline{h}_{i,N}~\mathrm{d}\mu\Bigg|=\,0.\end{split}$ (4.29)
@@
$\lim_{N\to\infty}\Bigg|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\tilde{a}_{N}(n)\int\prod_{i=1}^{2k}T^{[\tilde{f}_{i,m}(n)]}\tilde{h}_{i,N}~\mathrm{d}\mu\Bigg|=\,0.$ (4.30)
@@
[…]
5.  Proof of Theorem B
In this section we give the proof of Theorem B. The first step is to use ˜4.2, which was proved in the previous section, together with the structure theory of Host-Kra to reduce Theorem B to the following special case:
@@ 5.  Proof of Theorem B
Theorem B holds when $(X,\mathcal{B},\mu,T)$ is an ergodic nilsystem.
A measure preserving system is called an $s$-step pro-nilsystem if it is (isomorphic in the category of measure preserving systems to) an inverse limit of $s$-step nilsystems. In other words, $(X,\mathcal{B},\mu,T)$ is an $s$-step pro-nilsystem if there exist $T$-invariant $\sigma$-algebras ${\mathcal{B}}_{1}\subset{\mathcal{B}_{2}}\subset\cdots\subset{\mathcal{B}}$ such that ${\mathcal{B}}=\bigcup{\mathcal{B}}_{n}$ and for all $n$, the system $(X,{\mathcal{B}}_{n},\mu,T)$ is isomorphic to an $s$-step nilsystem. The following result from [HK05b] gives the relation between the uniformity seminorms and pro-nilsystems.
Theorem 5.2 ([HK05b]).
@@
Proof of Theorem B.
Let $\mathcal{H}$ be a Hardy field, let $f_{1},\ldots,f_{k}\in\mathcal{H}$ be of polynomial growth. Let $W\in\mathcal{H}$ be compatible with $f_{1},\dots,f_{k}$, in the sense that Property (P) holds. Let $(X,\mathcal{B},\mu,T)$ be an invertible measure preserving and let $h_{1},\dots,h_{k}\in L^{\infty}(X)$ and $A\in\mathcal{B}$ with $\mu(A)>0$. Using a usual ergodic decomposition argument, we may assume without loss of generality that the system is ergodic.
If $\lim_{t\to\infty}f_{i}(t)$ is finite for some $f_{i}$, then neither condition of Theorem A holds so parts (iii) and (ii) hold vacuously. As every $f_{i}\in\mathcal{H}$ is eventually monotone, this also implies that $f_{i}(n)]$ is eventually constant, and hence part (i) holds if it holds when $f_{i}$ is removed from $\{f_{1},\dots,f_{k}\}$. Iterating this observation we may assume that $\lim_{t\to\infty}|f_{i}(t)|=\infty$ for every $i$. A similar argument shows that we can assume that $\lim_{t\to\infty}|f_{i}(t)-f_{j}(t)|=\infty$ for every $i\neq j$.
Let $s\in\mathbb{N}$ be given by ˜4.2 and let ${\mathcal{Z}}_{s}$ be given by ˜5.2. In view of Theorems 4.2 and 5.2 we can assume that all the functions $h_{i}$ are measurable in ${\mathcal{Z}}_{s}$, which is equivalent to assuming that $(X,\mathcal{B},\mu,T)$ is itself an $s$-step pro-nilsystem. Therefore there exists a sequence of $\sigma$-algebras $\mathcal{B}_{1}\subset\mathcal{B}_{2}\subset\cdots\subset\mathcal{B}$ such that for each $\ell\in\mathbb{N}$ the system $(X,\mathcal{B}_{\ell},\mu,T)$ is an ergodic nilsystem.
It now follows from ˜5.1 that the conclusions (i) and (ii) of Theorem B hold after replacing each $h_{i}$ with $\mathbb{E}[h_{i}\mid\mathcal{B}_{\ell}]$, and that
$\lim_{N\to\infty}\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\int_{X}H_{\ell}\cdot T^{[f_{1}(n)]}H_{\ell}\cdots T^{[f_{k}(n)]}H_{\ell}\,>\,0$
where $H_{\ell}=\mathbb{E}[1_{A}\mid\mathcal{B}_{\ell}]$ and $\mathbb{E}[\cdot\mid\mathcal{B}_{\ell}]$ denotes the conditional expectation on the $\sigma$-subalgebra $\mathcal{B}_{\ell}$. Since $\mathbb{E}[h_{i}\mid\mathcal{B}_{\ell}]\to h_{i}$ and $H_{\ell}\to 1_{A}$ in $L^{p}$ for every $p<\infty$ as $\ell\to\infty$, we conclude that the conclusions of Theorem B also hold for $h_{1},\dots,h_{k}$ and $A$. ∎
The rest of this section is devoted to proving ˜5.1.
Outline of the proof of ˜5.1:
The first step in the proof of ˜5.1 is a reduction to the case when the nilmanifold $X$ is of the form $X=G/\Gamma$ for a connected and simply connected nilpotent Lie group $G$. These conditions on $G$ ensure that the nilsystem is embedable into a flow (i.e. a $\mathbb{R}$-action) on the same nilmanifold. This reduction is achieved at the price of losing ergodicity of the $\mathbb{Z}$-action; nevertheless, the $\mathbb{R}$-flow is ergodic. Then another reduction is performed, to avoid local obstructions.
These reductions allow us to reformulate ˜5.1 as a statement about uniform distribution of certain sequences in a nilmanifold, which we prove by drawing from recent results obtained in [Ric23].
@@ 5.1.  Reducing to connected and simply connected groups
The first step in the proof of ˜5.1 is essentially a reduction to the case when $X$ is a nilmanifold $G/\Gamma$ where the corresponding nilpotent Lie group $G$ is connected and simply connected. This is a somewhat technical and by now standard reduction but which is crucial in order to deal with the rounding function and to use the equidistribution results from [Ric23]. It turns out that the additional structure of nilsystems allows us to upgrade $L^{2}$ convergence into pointwise convergence. We now state this case of ˜5.1 as a separate theorem.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property (P). Then for any nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, where $X=G/\Gamma$ and $G$ is connected and simply connected, we have:
  1. (i)
@@
  2. (ii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (1) of Theorem A then for any $h_{1},\ldots,h_{k}\in C(X)$ and any $x\in X$
@@
  3. (iii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (2) of Theorem A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
    $\lim_{N\to\infty}\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\mu\Big(A\cap T^{-[f_{1}(n)]}A\cap\ldots\cap T^{-[f_{k}(n)]}A\Big)\,>\,0.$
Proof of ˜5.1 assuming ˜5.3.
Let $(X,\mathcal{B}_{X},\mu_{X},T)$ be an ergodic nilsystem. In view of ˜2.1 we can represent $X=G/\Gamma$ as a closed subsystem of another nilsystem $(Z,\mathcal{B}_{Z},\mu_{Z},S)$ where $Z=\hat{G}/\hat{\Gamma}$ for a connected and simply connected nilpotent Lie group $\hat{G}$ which contains $G$ as a closed and rational subgroup. Letting $\tilde{G}$ be the connected component of the identity in $G$ and letting $a\in G$ be such that $T$ corresponds to left multiplication by $a$, we also have that $G=\overline{a^{\mathbb{Z}}\tilde{G}}$ and $\hat{G}=\overline{a^{\mathbb{R}}\tilde{G}}$. If $G=\hat{G}$ then $G$ is connected and simply connected and the desired conclusion follows directly from ˜5.3. From now on suppose that $G\neq\hat{G}$.
To establish parts (i) and (ii) of Theorem B for the nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, it suffices to consider continuous functions $h_{i}$, for they are dense in $L^{2}$. By the Tietze-Urysohn extension theorem, any continuous function $h\in C(X)$ can be extended to a continuous function on $Z$. The conclusions now follow directly from parts (i) and (ii) of ˜5.3.
Before proving part (iii) of Theorem B for the nilsystem $(X,\mathcal{B}_{X},\mu_{X},T)$, we need some facts about the relation between $X$ and $Z$. For every $y\in X$ the set $D:=\{t\in\mathbb{R}:a^{t}y\in X\}=\{t\in\mathbb{R}:a^{t}\in G\}$ forms a closed subgroup of $\mathbb{R}$. We can’t have $D=\mathbb{R}$ for this would imply $G=\hat{G}$, so $D$ must be discrete. Notice that $D$ is also independent of $y$. Let $c=\min\{t>0:t\in D\}$. Let $R:=S^{c}:X\to X$. Since $1\in D$, it follows that $c=\tfrac{1}{m}$ for some $m\in\mathbb{N}$ and hence $R^{m}=T$ .
@@
Now, fix $A\subset X$ with $\mu_{X}(A)>0$. Let $B=\bigcup_{t\in[0,c)}S^{t}A$. It follows that $S^{-t}B\cap X=A$ for every $t\in[0,c)$ and hence $\mu_{Z}(B)=m\int_{0}^{c}\mu_{X}(a^{-t}B)~\mathrm{d}t=\mu_{X}(A)>0$. In view of part (iii) of ˜5.3 (and the choice of $W$) applied to $(Z,\mathcal{B}_{Z},\mu_{Z},S)$ we deduce that
@@
as long as the functions $f_{i}$ satisfy either condition (2) or (1) of Theorem A. Since $X$ is $T$ invariant, for any $n_{1},\dots,n_{k}\in\mathbb{Z}$ and any $t\in[0,c)$ we have